OpenStudy (anonymous):

A golfer hits a ball. The ball leaves the club with an initial upward velocity of 40 meters per second. If the ball is 0.25 meters off the ground when it leaves the golf club, about how many seconds will it take for the ball to hit the ground?

OpenStudy (anonymous):

Let's get some formula down x=X0+vt-1/2 g t^2

OpenStudy (amistre64):

h(t) = -meteric gravity units(t^2) +40t +.25

OpenStudy (amistre64):

4.6 or 4.9; cant recall the metric

OpenStudy (anonymous):

X0- initial height v- initial velocity g- gravity (9.8)

OpenStudy (amistre64):

h(t) = -4.9t^2 +40t +.25 i believe is good

OpenStudy (anonymous):

Set that equal to 0 and solve for t

OpenStudy (amistre64):

I get abt 8.1695..... seconds

OpenStudy (amistre64):

we can plug those into the quadratic formula to get an exact answer; \[\frac{-40 +\sqrt{40^2-4(-4.9)(.25)}}{2(-4.9)}\]